OpenStudy (anonymous):

simplify cubed root 108 please help

OpenStudy (anonymous):

factor 108

OpenStudy (anonymous):

u can do the factorization of 108 in prime numbers?

OpenStudy (anonymous):

ex 16=2*2*2*2 do the same for 108

OpenStudy (anonymous):

1st factor 108 into primes...did u do that?

OpenStudy (anonymous):

done

OpenStudy (anonymous):

suppose u have to find cube root of 8 cube root of 8= 8^(1/3)= (2*2*2)^(1/3) =(2^3)^(1/3)= 2

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):

108=2*2*3*3*3= 2^2 * 3^3

OpenStudy (anonymous):

so (108)^(1/3) = { 2^2 * 3^3} ^(1/3) = 2^(2/3) *3

OpenStudy (anonymous):

cube root 0f 108= 4^(1/3) * 3

OpenStudy (anonymous):

express the following surds in its simplest form of cube root of 108

OpenStudy (anonymous):

108^(1/3) = 4^(1/3) * 3

OpenStudy (anonymous):

or 108^(1/3) = 2^(2/3) * 3

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

why...wat wrong?

OpenStudy (anonymous):

r u 9 th std ? boy

OpenStudy (anonymous):

did u get the answer of ur question?

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

plz give the explaination if u know

OpenStudy (anonymous):

108=2*2*3*3*3= 2^2 * 3^3 so (108)^(1/3) = { 2^2 * 3^3} ^(1/3) = 2^(2/3) *3 did u get this?

OpenStudy (anonymous):

u know that power gets multiplied so(a^p)^(1/q) = (a) ^(p/q)

OpenStudy (anonymous):

and (ab)^n= (a)^n * (b) ^n

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

108=2*2*3*3*3= 2^2 * 3^3 so (108)^(1/3) = { 2^2 * 3^3} ^(1/3) = 2^(2/3) *3 did u get this?

OpenStudy (anonymous):

no.....

OpenStudy (anonymous):

ok 1st factor 108 into primes so start dividing 108 by smallest prime so 108=2*54 now divide 54 by 2 so 108=2*2*27 ............ 108= 2*2* 3*3*3

OpenStudy (anonymous):